OpenStudy (anonymous):

Allen’s boat used 6 gallons of gas in 5 hours. At this rate, how many gallons of gas will the boat use in 12 hours? a. 10 gallons c. 9.6 gallons b. 14 gallons d. 14.4 gallons

OpenStudy (anonymous):

set up a proportion... 6/5 = x/12 (6 gallons in 5 hrs = x gallons in 12 hrs) cross multiply (5)(x) = (6)(12) 5x = 72 -- divide both sides by 5 x = 72/5 x = 14 2/5 gallons or 14.4 gallons any questions ?

OpenStudy (anonymous):

no thanks!